OpenStudy (anonymous):

The width of a rectangle is 2 inches less than it's length, and the area is 35 square inches. What are the length and width of the rectangle?

myininaya (myininaya):

w=width l=length w=l-2 Area of rectangle =A=wl=(l-2)l=l^2-2l but the area is A=35 so l^2-2l=35 l^2-2l-35=0 so we need to solve this for l -7*5=-35 -7+5=-2 so l^2-2l-35=(l-7)(l+5)=0 l=7, l=-5 but it makes no sense for the length to be a negative so l=7 and if l=7, then w=7-2=5

OpenStudy (anonymous):

Thank you so much:D

myininaya (myininaya):

np